\begin{thebibliography}{xx}

\bibitem{didohand} A.~Di~Domenico (ed.), {\it Handbook on Neutral Kaon Interferometry at a $\phi$-factory}, {\em Frascati Phys. Ser.} {\bf 43}
INFN-LNF, Frascati, (2007); \\ available online at \\ http://www.lnf.infn.it/sis/frascatiseries/Volume43/volume43.pdf~.

\bibitem{ktviol} J.~Bernabeu, A.~Di~Domenico, P.~Villanueva-Perez,
%{\it Direct test of time reversal symmetry in the entangled neutral kaon
%system at a $\phi$-factory}, 
Nucl.\ Phys.\ B\ {\bf 868}, 102
%--119
 (2013).

\bibitem{kcptviol} J.~Bernabeu,  A.~Di~Domenico, P.~Villanueva-Perez,
%{\it Probing CPT in transitions with entangled neutral kaons},
J.\ High\ Energy\ Phys.\ {\bf 10}, 139 (2015).


\bibitem{kloeqm2} KLOE-2 Collaboration, D. Babusci \etal, 
%{\it Precision tests of quantum mechanics and CPT symmetry with entangled neutral kaons at KLOE}, 
J.\ High\ Energy\ Phys.\ {\bf 04}, 059 (2022).

\bibitem{future} J.~Bernabeu, A.~Di~Domenico, 
%{\it Can future observation of the living partner post-tag the past decayed state in entangled neutral K mesons?},
Phys.\ Rev\ D\ {\bf 105}, 116004 (2022).

\bibitem{didoactapp} A.~Di~Domenico, 
%{\it Testing discrete symmetries in transitions with entangled neutral kaons}, 
Acta Phys. Polonica B {\bf 48} (2017) 1919.
%-1924

%\bibitem{pdg} 
%P.A. Zyla et al. (Particle Data Group), The Review of Particle Physics (2020), Prog. Theor. Exp. Phys. 2020, 083C01 (2020).

\bibitem{kloetcpcpt} KLOE-2 Collaboration, D. Babusci \etal, in preparation.

\bibitem{datatables}
{\it Data Tables for Lorentz and CPT Violation,}
V.A.\ Kosteleck\'y and N.\ Russell,
2022 edition,
arXiv:0801.0287v15.

\bibitem{kost1} {
V.~A.~Kosteleck\'y,  
%Sensitivity of CPT Tests with Neutral Mesons. 
{ Phys. Rev. Lett.} {\bf 80}, 1818
%--1821.
(1998).
}

\bibitem{kost2} {
V.~A.~Kosteleck\'y, 
%Signals for CPT and Lorentz violation in neutral-meson oscillations. 
{Phys. Rev. D} {\bf 61}, 016002 (1999).
}

\bibitem{kost3} {
V.~A.~Kosteleck\'y, 
%CPT, T, and Lorentz violation in neutral-meson oscillations. 
{Phys. Rev. D} {\bf 64}, 076001 (2001).
}

\bibitem{kost4} {
B.~R.~Edwards, V.~A.~Kosteleck\'y, 
%Searching for CPT violation with neutral-meson oscillations
{Phys.Lett. B} {\bf 795}, 620
%--626.
(2019).
}

\bibitem{agnes} \'A.~Roberts, 
%Testing CPT symmetry with correlated neutral mesons, 
Phys. Rev. D {\bf 96}, 116015 (2017).

\bibitem{agnesralf} R.~Lehnert, \'A.~Roberts, in preparation.

\bibitem{kloecpt2013} 
KLOE-2 Collaboration, D.~Babusci \etal,
%Test of CPT and Lorentz symmetry in entangled neutral kaons with the KLOE experiment,
Phys.\ Lett.\ B {\bf 730}, 89
%--94 
(2014).

%
%
%\bibitem{colloquium} J.~Bernabeu and F.~Martinez-Vidal, 
%{\it Colloquium: Time-reversal violation with quantum-entangled B mesons}, 
%Rev.~Mod.~Phys. {\bf 87}, 165 (2015)
%
%\bibitem{tviol} J. Bernabeu, A. Di Domenico, P. Villanueva-Perez, 
%{\it Direct test of time reversal symmetry in the entangled neutral kaon system at a $\phi$-factory}, 
%Nucl. Phys. B. {\bf 868}, 102 (2013)
%
%
%\bibitem{cpttran} J. Bernabeu, A. Di Domenico, P. Villanueva-Perez, 
%{\it Probing CPT in transitions with entangled neutral kaons}, 
%JHEP {\bf 10}, 139 (2015)
%
%\bibitem{ref:HandbookAD} A. Di Domenico (Editor) , {\it Handbook on Neutral Kaon Interferometry at a $\phi$-factory}, Frascati Phys. Ser. {\bf 43} (2007)
%
%\bibitem{kloe2epjc} G.~Amelino-Camelia et al., 
%{\it Physics with the KLOE-2 experiment at the upgraded DA$\Phi$NE}, 
%Eur. Phys. J. C {\bf 68}, 619 (2010)
%
%\bibitem{gajos} A.~Gajos, these proceedings.
%
%
%
%\bibitem{randomphoton}
%J.\ Lipa, J.A.\ Nissen, S.\ Wang, D.A.\ Stricker, and D.\ Avaloff,
%Phys.\ Rev.\ Lett.\ {\bf 90}, 060403 (2003);
%H.\ M\"uller, S.\ Herrmann, C.\ Braxmaier,
%S.\ Schiller, and A.\ Peters, 
%Phys.\ Rev.\ Lett.\ {\bf 91}, 020401 (2003);
%P.\ Wolf, M.E.\ Tobar, S.\ Bize, A.\ Clairon,
%A.N.\ Luiten, and G.\ Santarelli,
%Gen.\ Rel.\ Grav.\ {\bf 36}, 2351 (2004).
%
%\bibitem{randomnu}
%MINOS Collaboration,
%P.\ Adamson \etal,
%Phys.\ Rev.\ Lett.\ {\bf 101}, 151601 (2008);
%LSND Collaboration,
%L.B.\ Auerbach \etal, 
%Phys.\ Rev.\ D {\bf 72}, 076004 (2005);
%J.S.\ D\'\i az \etal,
%Phys.\ Rev.\ D {\bf 80}, 076007 (2009).
%
%\bibitem{randombook}
%F.\ Boehm and P.\ Vogel,
%{\it Physics of Massive Neutrinos},
%Cambridge University Press, Cambridge, 1987.

\end{thebibliography}

\documentclass{ws-procs9x6-cpt22}
\begin{document}
\newcommand{\refeq}[1]{(\ref{#1})}
\def\etal {{\it et al.}}
%
\newcommand{\pn}{\ensuremath{\mathrm{p}}}
\newcommand{\pnb}{\ensuremath{\mathrm{\bar{p}}}}
\newcommand{\ksn}{\ensuremath{\mathrm{K_S}}}
\newcommand{\ksln}{\ensuremath{\mathrm{K_{S,L}}}}
\newcommand{\kln}{\ensuremath{\mathrm{K_L}}}
\newcommand{\ksnr}{\ensuremath{\mathrm{K_S^R}}}
\newcommand{\klnr}{\ensuremath{\mathrm{K_L^R}}}
\newcommand{\kkn}{\ensuremath{\mathrm{K}}}
\newcommand{\kknt}{\ensuremath{\mathrm{\widetilde K}}}
\newcommand{\kn}{\ensuremath{\mathrm{K^0}}}
\newcommand{\knb}{\ensuremath{\mathrm{\bar{K}^0}}}
\newcommand{\mn}{\ensuremath{\mathrm{M^0}}}
\newcommand{\mnb}{\ensuremath{\mathrm{\bar{M}^0}}}
\newcommand{\kpp}{\ensuremath{\mathrm{K_+}}}
\newcommand{\kppnn}{\ensuremath{\mathrm{K_{\pm}}}}
\newcommand{\knn}{\ensuremath{\mathrm{K_-}}}
\newcommand{\kpport}{\ensuremath{\mathrm{K_+^{\perp}}}}
\newcommand{\knnpport}{\ensuremath{\mathrm{K_{\mp}^{\perp}}}}
\newcommand{\knnort}{\ensuremath{\mathrm{K_-^{\perp}}}}
\newcommand{\knnmp}{\ensuremath{\mathrm{K_{\mp}}}}
\newcommand{\mpp}{\ensuremath{\mathrm{M_+}}}
\newcommand{\mnn}{\ensuremath{\mathrm{M_-}}}
\newcommand{\kppppm}{\ensuremath{\mathrm{\widetilde K_{\pm}}}}
\newcommand{\kppp}{\ensuremath{\mathrm{\widetilde K_+}}}
\newcommand{\knnp}{\ensuremath{\mathrm{\widetilde K_-}}}
\newcommand{\knt}{\ensuremath{\mathrm{\widetilde K_0}}}
\newcommand{\knbt}{\ensuremath{\mathrm{\widetilde K_{\bar{0}}}}}
\newcommand{\kntt}{\ensuremath{\mathrm{ K_0}}}
\newcommand{\knbtt}{\ensuremath{\mathrm{ K_{\bar{0}}}}}
\newcommand{\bn}{\ensuremath{\mathrm{B^0}}}
\newcommand{\bbn}{\ensuremath{\mathrm{B}}}
\newcommand{\bbnd}{\ensuremath{\mathrm{B_d}}}
\newcommand{\bnb}{\ensuremath{\mathrm{\bar{B}^0}}}
\newcommand{\kknx}{\ensuremath{\mathrm{K_X}}}
\newcommand{\kkny}{\ensuremath{\mathrm{K_Y}}}
\newcommand{\kkone}{\ensuremath{\mathrm{K_1}}}
\newcommand{\kktwo}{\ensuremath{\mathrm{K_2}}}
\newcommand{\kknxb}{\ensuremath{\mathrm{\bar{K}_X}}}
\newcommand{\kknyb}{\ensuremath{\mathrm{\bar{K}_Y}}}
\newcommand\trule{\rule{0pt}{2.6ex}}

\def\CP       {\ensuremath{\mathrm{CP}}}
\def\CPT       {\ensuremath{\mathrm{CPT}}}
\def\C       {\ensuremath{\mathrm{C}}}
\def\P       {\ensuremath{\mathrm{P}}}
\def\T       {\ensuremath{\mathrm{T}}}
\def\SS    {\ensuremath{\mathrm{S}}}
\def\K       {\ensuremath{K}}
\def\B       {\ensuremath{B}}

\title{Entanglement, CPT and neutral kaons}

\author{Antonio Di Domenico}

\address{Dipartimento di Fisica, Sapienza Universit\`a di Roma and
INFN Sezione di Roma \\
P.\ le A.\ Moro 2, I-00185, Rome, Italy}
%

\begin{abstract}
Direct tests of \T, \CP, \CPT\ symmetries in transitions
using entangled neutral kaons produced at a $\phi$-factory
are briefly reviewed.
%
\end{abstract}

\bodymatter

\section{Introduction}
%
Neutral kaons produced at a $\phi$-factory combine their 
peculiar
flavour oscillations, charge-parity (\CP) and time-reversal (\T) violation,
into an Einstein-Podolsky-Rosen (EPR) entangled system, revealing
surprising features.\cite{didohand,ktviol,kcptviol,kloeqm2,future}
Here 
the possibility of exploiting 
their entanglement 
to make direct tests of \T, \CP, \CPT\ symmetries
is briefly reviewed.

\section{Direct test of 
discrete symmetries with neutral kaons}
\par
In order 
to implement direct tests of \T, \CP, \CPT\  symmetries in transitions, the preparation of the initial kaon state
is performed exploiting the 
entanglement of $\kn\knb$ pairs produced at a $\phi$-factory.\cite{didohand,ktviol,kcptviol}
In this way, taken a kaon transition process as a reference,
the exchange of 
{\it in} and {\it out} states 
required for a genuine 
test 
involving an anti-unitary transformation implied by time-reversal,
can be easily implemented.
In fact,
the initial kaon pair produced in $\phi\rightarrow \kn\knb$ decays
can be rewritten in terms of any pair of orthogonal states:
\begin{eqnarray}
  |i \rangle   =  \frac{1}{\sqrt{2}} \{ |\kn \rangle |\knb \rangle - 
 |\knb \rangle |\kn \rangle
\} 
\label{eq:state1}
   =  \frac{1}{\sqrt{2}} \{ |\kpp \rangle |\knn \rangle - 
 |\knn \rangle |\kpp \rangle
\label{eq:state3}
\}~.
\end{eqnarray}  
Here 
the states $|\knn\rangle$, $|\kpp\rangle$ are defined as 
the states which cannot decay into pure $\CP=\pm1$ final states,
$\pi\pi$ 
or $3\pi^0$, respectively.\cite{ktviol,kcptviol}
%
%
%
The condition of orthogonality $\langle\knn|\kpp\rangle=0$, 
corresponds to assume negligible direct \CP\ (or \CPT) violation contributions, while 
the $\Delta S=\Delta Q$ rule is also assumed, so that the two flavor orthogonal eigenstates $|\kn\rangle$ and $|\knb\rangle$ are identified by the charge of the lepton in semileptonic decays.
%
\par
Thus, exploiting the perfect anticorrelation of the states implied 
by Eq.~\refeq{eq:state3},
 it is possible to have a 
\textquotedblleft flavor-tag\textquotedblright or a 
 \textquotedblleft \CP-tag\textquotedblright,
i.e.~to infer the flavor (\kn or \knb) or the \CP\ (\kpp or \knn) state
of the still alive kaon by observing a specific flavor decay ($\pi^+\ell^-\nu$
or  $\pi^-\ell^+\bar{\nu}$) or CP decay ($\pi\pi$ or $3\pi^0$) of the other (and first decaying) kaon in the pair.
Then the decay of the surviving kaon into a semileptonic ($\ell^+$ or $\ell^-$), $\pi\pi$ or $3\pi^0$ final state, filter the kaon final state as a flavor or \CP\ 
state.
%
\par
In this way one can identify 
a reference transition (e.g. $\kn \to \knn$)
and its symmetry conjugate (e.g. the \CPT-conjugated $\knn \to \knb$),
and directly compare them
through
the 
corresponding 
ratios of probabilities.
The observable ratios 
for the various symmetry tests can be defined as follows:\cite{didoactapp}
%
%
\begin{eqnarray}
R_{2,\T}^{\rm{exp}}(\Delta t) &\equiv&
\frac{  I(\ell^-,3\pi^0;\Delta t)}
{ I(\pi\pi,\ell^+;\Delta t)} \cdot \frac{1}{D_{\CPT}}
\nonumber\\
&=&\left(1-4\Re \epsilon +4\Re x_+ + 4 \Re y \right) 
\nonumber\\
& & \times 
\left| 1+ 
\left(
2\epsilon
+\epsilon^{\prime}_{3\pi^0} +\epsilon^{\prime}_{\pi\pi}
\right)
e^{-i(\lambda_S-\lambda_L)\Delta t} \right|^2 
~,
\label{ratio2texp}
\end{eqnarray}
\begin{eqnarray}
R_{4,\T}^{\rm{exp}}(\Delta t) &\equiv&
\frac{  I(\ell^+,3\pi^0;\Delta t)}
{ I(\pi\pi,\ell^-;\Delta t)}   \cdot \frac{1}{D_{\CPT}}
\nonumber\\
&=&\left(1+4\Re \epsilon +4\Re x_+ - 4 \Re y \right) 
\nonumber\\
& & \times \left| 1-
\left(
2\epsilon
+\epsilon^{\prime}_{3\pi^0} +\epsilon^{\prime}_{\pi\pi}
\right)
e^{-i(\lambda_S-\lambda_L)\Delta t} \right|^2
~,
\label{ratio4texp}
\end{eqnarray}
\begin{eqnarray}
R_{2,\CP}^{\rm{exp}}(\Delta t) &\equiv&
\frac{  I(\ell^-,3\pi^0;\Delta t)}
{ I(\ell^+,3\pi^0;\Delta t)}
\nonumber\\
&=&\left(1-4\Re \epsilon_S 
-4\Re x_- + 4 \Re y \right) 
\nonumber\\
& & \times \left| 1+ 
2\left(
\epsilon_S 
+\epsilon^{\prime}_{3\pi^0} 
\right)
e^{-i(\lambda_S-\lambda_L)\Delta t} \right|^2 ~,
\label{ratio2cpexp}
\end{eqnarray}
\begin{eqnarray}
R_{4,\CP}^{\rm{exp}}(\Delta t) &\equiv&
\frac{  I(\pi\pi,\ell^+;\Delta t)}
{ I(\pi\pi,\ell^-;\Delta t)}   
\nonumber\\
&=&\left(1+4\Re \epsilon_L 
-4\Re x_- - 4 \Re y \right) 
\nonumber\\
& & \times \left| 1-
2\left(
\epsilon_L
+\epsilon^{\prime}_{\pi\pi}
\right)
e^{-i(\lambda_S-\lambda_L)\Delta t} \right|^2 ~,
\label{ratio4cpexp}
\end{eqnarray}
\begin{eqnarray}
R_{2,\CPT}^{\rm{exp}}(\Delta t) &\equiv&
\frac{  I(\ell^-,3\pi^0;\Delta t)}
{ I(\pi\pi,\ell^-;\Delta t)} \cdot \frac{1}{D_{\CPT}}
\nonumber\\
&=& \left(1-4\Re \delta+4\Re x_+ - 4 \Re x_-\right) 
\nonumber\\
& & \times 
\left| 1+ 
\left(
2\delta 
+\epsilon^{\prime}_{3\pi^0} -\epsilon^{\prime}_{\pi\pi}
\right)
e^{-i(\lambda_S-\lambda_L)\Delta t} \right|^2
~,
\label{ratio2cptexp}
\end{eqnarray}
\begin{eqnarray}
R_{4,\CPT}^{\rm{exp}}(\Delta t) &\equiv&
\frac{  I(\ell^+,3\pi^0;\Delta t)}
{ I(\pi\pi,\ell^+;\Delta t)} \cdot \frac{1}{D_{\CPT}}
\nonumber\\
&=& \left(1+4\Re \delta+4\Re x_+ + 4 \Re x_-\right) 
\nonumber\\ 
& & \times 
\left| 1-
\left(
2\delta 
+\epsilon^{\prime}_{3\pi^0} -\epsilon^{\prime}_{\pi\pi}
\right)
%
%
%
e^{-i(\lambda_S-\lambda_L)\Delta t} \right|^2
~,
\label{ratio4cptexp}
\end{eqnarray}
where
$I(f_1,f_2;\Delta t)$
are the double decay rates
into 
decay products
$f_1$ 
and $f_2$ 
as a function of the difference of kaon decay times $\Delta t$,\cite{didohand,ktviol,kcptviol} with
$f_1$ occurring before $f_2$ decay for $\Delta t >0$, and viceversa for $\Delta t<0$.
$D_{\CPT}$ is a constant factor that can be determined from measurable branching fractions and lifetimes of \ksln\ states~\cite{kcptviol}.
For $\Delta t=0$ one has by construction, within our assumptions, $R_{i,\SS}^{\rm{exp}}(0)=1$ (with $\SS=\T,\CP$, or $\CPT$, and $i=2,4$), and
the measurement of any deviation from the prediction $R_{i,\SS}^{\rm{exp}}(\Delta t)=1$
imposed by the symmetry invariance 
is a 
direct signal of the symmetry violation built in the time 
evolution of the system.
The following double ratios independent of the factor $D_{\CPT}$ can also be defined:
\begin{eqnarray}
\label{eq:dratiot}
DR_{\T,\CP} (\Delta t) &\equiv & \frac{R_{2,\T}^{\rm{exp}}(\Delta t)}{ R_{4,\T}^{\rm{exp}}(\Delta t)} \equiv \frac{R_{2,\CP}^{\rm{exp}}(\Delta t)}{R_{4,\CP}^{\rm{exp}}(\Delta t)} 
\nonumber\\
&=& \left(1-8\Re \epsilon+8\Re y \right) 
\nonumber\\ 
& & \times 
\left| 1+2
\left(
2\epsilon
+\epsilon^{\prime}_{3\pi^0} +\epsilon^{\prime}_{\pi\pi}
\right)
%
%
%
e^{-i(\lambda_S-\lambda_L)\Delta t} \right|^2
~,
\\
\label{eq:dratiocpt}
DR_{\CPT} (\Delta t) & \equiv & \frac{R_{2,\CPT}^{\rm{exp}}(\Delta t)}{R_{4,\CPT}^{\rm{exp}}(\Delta t)} 
\nonumber\\
&=& \left(1-8\Re \delta-8\Re x_-\right) 
\nonumber\\ 
& & \times 
\left| 1+2
\left(
2\delta 
+\epsilon^{\prime}_{3\pi^0} -\epsilon^{\prime}_{\pi\pi}
\right)
%
%
%
e^{-i(\lambda_S-\lambda_L)\Delta t} \right|^2
~.
\end{eqnarray}  
The r.h.s. of Eqs.~\refeq{ratio2texp}-\refeq{eq:dratiocpt} 
is evaluated to first order in small parameters
and for not too large negative $\Delta t$;
$\epsilon$ and $\delta$ are the usual \T\ and \CPT\ violation parameters in the neutral
kaon mixing, respectively, and $\epsilon_{S,L}=\epsilon\pm\delta$ the \CP\ impurities in the 
physical states \ksn\ and \kln;
the small parameter $y$ describes a possible \CPT\ violation in the $\Delta S = \Delta Q$ semileptonic decay amplitudes, while
$x_+$ and $x_-$ describe $\Delta S \neq \Delta Q$ semileptonic decay amplitudes
with \CPT\ invariance and \CPT\ violation, respectively.
Therefore the r.h.s. of Eqs.~\refeq{ratio2texp}-\refeq{eq:dratiocpt} 
shows
the effect of
symmetry 
violations only in the
the effective Hamiltonian description of the neutral kaon system according to the
Weisskopf-Wigner approximation,
without the presence of other possible sources of symmetry violations.
The small spurious effects due to the release of our assumptions are also shown, including possible $\Delta S=\Delta Q$ rule violations ($x_+,x_-\neq0$)
and/or direct \CP\ and/or \CPT\ violation effects ($\epsilon^{\prime}_{3\pi^0},\epsilon^{\prime}_{\pi\pi},y\neq0$). In particular
the $\epsilon^{\prime}$ effects are fully negligible in the asymptotic region $\Delta t \gg \tau_S$.
\par
The KLOE-2 collaboration has recently analysed a data sample corresponding to an integrated luminosity $L=1.7~\hbox{fb}^{-1}$ collected 
at the DA$\Phi$NE $\phi$-factory, and measured all eight observables defined in Eqs.\ \refeq{ratio2texp}-\refeq{eq:dratiocpt} in the asymptotic region $\Delta t \gg \tau_S$ (with positive $\Delta t$).\cite{kloetcpcpt} These results constitute the first direct tests of \T\ and \CPT\ symmetries in kaon transitions.
The \T\ and \CPT\ observables have been measured with a precision of few percent, showing no evidence of symmetry violations, 
while \CP\ violation has been observed with more than $5\sigma$ significance with the observable ratio \refeq{ratio4cpexp}.
 \par
The double ratio $DR_{\CPT}$
in  Eq.\ \refeq{eq:dratiocpt} 
in the asymptotic regime 
appears one of 
the best observable for testing \CPT, free from approximations and model independent~\cite{kcptviol}.
Thus it seems best suited to extend the \CPT\ tests to
the general framework of the Standard Model Extension (SME) for \CPT\ and Lorentz symmetry breaking,\cite{datatables,kost1,kost2,kost3,agnes,kost4,agnesralf}
e.g. by studying the \CPT\ observables 
as a function of 4-momenta of the kaons.\cite{kloecpt2013}
\par Finally it is worth noting that the full exploitation of time correlations of the two entangled kaons, not only {\it from past to future},
as in the above mentioned direct \T, \CP, and \CPT\ tests, but also {\it from future to past}, as described in Ref.\ \refcite{future}
to post-tag special kaon states,
might lead to the identification of new observables suitable to further extend these symmetry tests.

\section*{Acknowledgments}
The author 
would like to warmly thank R.~Lehnert for the invitation to this very interesting edition of the meeting.

\end{document}